### v1

# Dynamic Language Group-Based MoE: Enhancing Efficiency and Flexibility for Code-Switching Speech Recognition

Abstract
The Mixture of Experts (MoE) approach is ideally suited for tackling multilingual and code-switching (CS) challenges due to its multi-expert architecture. This work introduces the DLG-MoE, which is optimized for bilingual and CS scenarios. Our novel Dynamic Language Group-based MoE layer features a language router with shared weights for explicit language modeling, while independent unsupervised routers within the language group handle attributes beyond language. This structure not only enhances expert extension capabilities but also supports dynamic top-$k$ training, allowing for flexible inference across various top-$k$ values and improving overall performance. The model requires no pre-training and supports streaming recognition, achieving state-of-the-art (SOTA) results with unmatched flexibility compared to other methods. The Code will be released.
Index Terms—  Mixture of experts, code-switching, automatic speech recognition
1 Introduction
Thanks to its ability to scale parameters with minimal impact on inference cost, the Mixture of Experts (MoE) approach has achieved exceptional results in natural language processing and computer vision [1, 2, 3]. The inherent multi-expert structure of MoE provides a natural advantage for multilingual tasks. Numerous studies [4, 5, 6] have demonstrated its effectiveness in multilingual automatic speech recognition (ASR), highlighting MoE’s strong performance in this domain. However, these architectures still struggle to effectively handle CS scenarios.
CS scenarios often involve using different languages within utterances, leading to significant challenges in model development due to language confusion. In response, Bi-encoder architectures [7, 8, 9, 10] use separate encoders for each language and employ language-specific masking strategies [8, 9, 11] to enhance individual language processing. However, this approach fails to leverage shared representations across languages, resulting in the loss of cross-lingual contextual information and increased computational complexity. LSR-MoE [12] and LR-MoE [13] mitigate these issues by incorporating two experts within the MoE layer, each dedicated to handling Chinese or English separately, and activate only one expert at a time during training and inference for computational efficiency. Similarly, Google [4] uses multiple experts in the MoE layer, activating two experts to manage multilingual ASR. U2++ MoE [14] highlights MoE’s effectiveness on large-scale Chinese datasets, providing both streaming and non-streaming capabilities within a single MoE and outperforming other models in this aspect.
In this paper, we present DLG-MoE, an efficient and flexible MoE structure. Addressing limitations of previous approaches, we designed language expert groups and a language router with weight sharing at the MoE layer to support flexible expansion. The language router explicitly learns language identification (LID), focusing on distributing frame-level representations to matching language groups based on LID. Within each language group, an unsupervised router specializes in modeling attributes beyond language, such as accents and different domains, distributing inputs to appropriate experts within the group. We also introduce a dynamic top-$k$ routing mechanism during training, allowing for the activation of a variable number of experts during inference. Our contributions are summarized as follows:
  • Our DLG-MoE achieves SOTA performance compared to existing MoE works [7, 12, 13] and also offers streaming recognition capabilities and greater scalability, supporting language-specific and different top-$k$ inference.
  • This paper further investigates the impact of different routing strategies and increasing the number of experts on model performance and also validates the approach on larger datasets and with larger parameters.
Fig. 1: (a) Overall Architecture, (b) DLG-MoE layer, (c) Dynamic Language Group.
2 Proposed Method
2.1 Overall Model Architecture
The overall architecture of our model is shown in Fig.1(a), comprising Conformer [15] Blocks and DLG-MoE Blocks. The Conformer Block is stacked from vanilla Conformer layers, while the DLG-MoE Block is formed by stacking DLG-MoE layers. Fig.1(b) shows the structure of the DLG-MoE layer, where the second FFN of the Conformer layer is replaced by a dynamic language group. The weight-shared language router (SLR) performs the LID task based on the intermediate representations $h_{inter}$ extracted by the Conformer block and uses the LID results to guide the selection of the appropriate language group.
2.2 Language Router
As illustrated in Fig.1(c), we expand MoE by introducing language expert groups, specifically the Chinese expert group (Zh-Group) and the English expert group (En-Group). Unlike other methods [14, 12, 4], our language router explicitly learns frame-level LID and is trained using an inter-loss function, as expressed by the following formula Equation1.
$\begin{split}\mathcal{L}_{\text{inter}}=&-\left(\log\mathrm{P}_{\mathrm{CTC}}\left(\mathcal{Y}_{\mathrm{LID}}\mid\mathrm{h}_{\text{inter}}\mathrm{W}_{\text{LID}}^{d\times\mathrm{L}+1}\right)\right.\\ &\left.+\log\mathrm{P}_{\mathrm{CTC}}\left(\mathcal{Y}_{\mathrm{ASR}}\mid\mathrm{h}_{\text{inter}}\mathrm{W}_{\text{ASR}}^{d\times V}\right)\right)\end{split}$ (1)
We enhance the SLR’s language routing capability through a multi-task training approach that integrates ASR and LID tasks. Here, $\mathcal{Y}_{\text{ASR}}$ denotes the ground truth for the ASR task, while $\mathcal{Y}_{\text{LID}}$ is derived by mapping the tokens in $\mathcal{Y}_{\text{ASR}}$ to their corresponding language IDs. The linear layers $\mathrm{W}_{\text{LID}}^{d\times\mathrm{L}+1}$ and $\mathrm{W}_{\text{ASR}}^{d\times V}$ represent the transformations, where $d$ is the embedding dimension, $V$ is the vocabulary size, and $L$+1 accounts for the number of languages plus the CTC blank token.
Considering the peaky distribution of CTC which classifies most frames as blank, we use a simple yet empirically effective routing strategy. For the logits from the SLR, which a shape of $\mathbb{R}^{T\times(L+1)}$, where $T$ represents the number of frames, we first remove the blank class to get $\mathbb{R}^{T\times L}$, then select the highest probability among the remaining categories as the language ID for every frame, resulting in a routing table of shape $\mathbb{R}^{T\times 1}$. This greedy strategy, which does not rely on other frames, is effective and beneficial to support streaming ASR. Thus, for the input $h_{pre}$ to the dynamic language group, it is dispatched as $h_{pre}^{zh}$ and $h_{pre}^{en}$ based on the routing information—specifically, the LID results—provided by the SLR, and then processed by the corresponding language expert groups.
2.3 Dynamic Language Group
As shown in Fig. 1(c), once the representations for the same language are obtained, the UnSup-Router within the language expert group focuses on modeling attributes beyond language. For each frame in $h_{pre}^{zh}$, the UnSup-Router selects $k$ experts from the total $n$ experts within the group to collaborate, as described below:
$h_{\text{out }}^{zh}=\sum_{i=1}^{k}G_{i}^{r}E_{i}$ (2)
$G^{r}=\operatorname{Softmax}\left(\operatorname{TopK}\left(h_{pre}^{zh}W_{R}^{d\times n},k\right)\right)$ (3)
Here, $G_{i}^{r}$ and $E_{i}$ denote the representations from the selected experts and their corresponding weights. In Equation 3, $W_{R}^{d\times n}$ is a linear layer that projects the representation $h_{pre}^{zh}$ onto the number of experts $n$ and selects $k$ experts for activation. It is important to note that different frames may be assigned to different sets of $k$ experts, which are then combined in a weighted manner to produce $h_{out}^{zh}$. Subsequently, the Zh-Group and En-Group extract specific language representations $h_{out}^{zh}$ and $h_{out}^{en}$, respectively, combine them to obtain $h_{out}$, and pass it to the next layer. To support various top-$k$ inferences, we employ a dynamic top-$k$ strategy within the expert groups, where the UnSup-Router randomly selects the values of top-k during training. It is crucial to emphasize that neither the language expert groups nor the language router require pre-training; the entire model can be initialized randomly and trained from scratch. All losses are jointly optimized, as illustrated below:
$\mathcal{L}_{\text{total }}=\lambda_{ctc}\mathcal{L}_{\text{ctc }}+(1-\lambda_{ctc})\mathcal{L}_{\text{att}}+\lambda_{\text{inter}}\mathcal{L}_{\text{inter}}$ (4)
3 Experiments and Analysis
3.1 Experiment setup
Our experiments are conducted on the ASRU-2019 Mandarin-English code-switching challenge dataset [16], along with a subset of 460 hours from Librispeech [17], consistent with [7, 12, 13]. For testing, we use Librispeech-clean (EN), ASRU-Man-test (ZH), and ASRU-CS-test (CS) datasets. Additionally, we have validated our model on an internal multi-accent, multi-domain dataset, which, combined with the ASRU-2019 Challenge dataset, totals 6,000 hours of training data.
We conduct our experiments on the WeNet [18] toolkit. For acoustic features, 80-dimensional log-mel filter banks (FBank) are extracted with a step size of 10ms and a window size of 25ms. All our experimental configurations used a 6-layer Transformer decoder. We set $d^{model}$=256, $n^{heads}$=4 and $d^{ffn}$=2048. Causal convolution was used during training to support streaming. We built Baseline-12 and Baseline-18 with 12 and 18 Conformer layers, respectively. All MoE models had 12 layers, with the first 6 being vanilla Conformer layers and the last 6 incorporating MoE layers. Dense-MoE-4e and Sparse-MoE-4e, which do not include language routing, each had 4 MoE experts. In contrast, DLG-MoE-4e, also with 4 experts, included language routing, allocating two experts per language group. To explore model capacity and different top-$k$ strategies, we created DLG-MoE-8e with 8 experts. For large-scale comparisons, we used $d^{model}$=512 and $n^{heads}$=8 for Baseline-18 and DLG-MoE-8e, while keeping other configurations the same. The hyperparameters $\lambda_{ctc}$ and $\lambda_{inter}$ were set to 0.3 and 0.1, respectively, with the Adam optimizer using $lr$=0.001 and $warmup_{s}teps$=25000. Detailed information can be found in the link provided  [^1] https://github.com/kaihuhuang/Language-Group
3.2 Results of different MoE structures
The results in Table 1 show that MoE structures ($S3$, $S4$) outperform $S2$ with the same number of parameters, indicating that widening the model with MoE might more effective than simply increasing he number of layers. This suggests that even when using only the UnSup-Router, experts can still implicitly learn to model different attributes, such as language or accent, leading to better results. $S4$, which activates only half the experts compared to $S3$, performs better, showing that selective activation of experts is more reasonable. Our proposed $S6$ shows a significant performance improvement over $S3$ and $S4$, proving the effectiveness of language expert groups. The language router’s explicit language differentiation ensures that our language groups focus on specific language representations without interference. In $S7$, without the guidance of an unsupervised router, the expert weights within the language group are assigned the same value, contrasting with the situation in experiment $S6$. In $S6$, the expert weights are dynamically learned through an unsupervised router, highlighting the significant impact of the learned weights in the UnSup-Router. DLG-MoE’s language router focuses on explicit language modeling, allowing the unsupervised routers within the language groups to model attributes beyond language specifically, such as accent or different domains, resulting in better performance. Our model achieves SOTA performance on the CS test set compared to previous MoE works. Specifically, it outperforms Biencoder-MoE(9.76%) [7] , LSR-MoE(10.30%) [12] , and FLR-MoE(9.7%) [13], all using the same training data, while also offering higher flexibility and customizability.
Table 1: Performance of different systems on the test sets. For EN, ZH and CS, we report the word error rate (WER), character error rate (CER), and mixed error rate (MER), respectively. We also assess the model’s computational complexity by evaluating the floating-point operations (FLOPs) for a 20-second input sample.
System Model Top-k Train Size Inference Size FLOPs↓ EN ZH CS
$S1$ Baseline-12 top1 51M 51M(top-1 decode) 24.8G 6.89 3.35 9.93
$S2$ Baseline-18 top1 72M 72M(top-1 decode) 36.2G 6.52 2.93 9.52
$S3$ Dense-MoE-4e top4 72M 72M(top-2 decode) 34.3G 6.09 2.73 9.36
$S4$ Sparse-MoE-4e top2 72M 60M(top-2 decode) 27.8G 5.99 2.70 9.38
$S5$ Sparse-MoE-4e dynamic 72M 60M(top-2 decode) 27.8G 6.12 2.89 9.62
$S6$ DLG-MoE-4e top2 72M 51M(top-1 decode) 25.0G 5.29 2.77 9.81
60M(top-2 decode) 27.8G 5.31 2.55 9.09
[…]
DLG-MoE-4e
(w/o Router)
top2 72M 60M(top-2 decode) 27.8G 5.44 2.67 9.23
[…]
60M(top-2 decode) 27.8G 5.42 2.52 9.10
$S9$ DLG-MoE-8e top1 103M 51M(top-1 decode) 25.0G 5.62 2.58 9.30
60M(top-2 decode) 27.8G 5.82 2.66 9.52
[…]
60M(top-2 decode) 27.8G 5.26 2.39 9.10
[…]
60M(top-2 decode) 27.8G 5.29 2.42 8.92
3.3 Analysis of expert capacity and top-k
[…]
Table 2: Results with large-scale datasets and parameters.
{tblr}
cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=3c, cell41 = r=2, cell42 = r=2, hline1,6 = -0.08em, hline2 = 4-6, hline3 = 1-3, hline3 = 4-60.03em, hline4 = -dashed, Model & Size Inf. Size ASRU-CS
MER CER WER
Baseline-18 173M 173M(top1) 6.65 4.50 24.25
DLG-MoE-8e 228M 100M(top1) 6.16 4.12 22.88
118M(top2) 6.11 4.08 22.78
3.4 Streaming Capability
Table 3: Streaming results for different chunk sizes.
Mode Specify Lang. EN ZH CS
NonStreaming 5.29 2.42 8.92
Streaming(640ms) 6.21 2.82 9.89
Streaming(320ms) 6.69 3.02 10.31
Streaming(320ms) 6.64 2.98 10.31
In this section, we tested the model’s streaming recognition ability with different chunk sizes. As illustrated in Table 3, our model supports inference with varying chunk sizes while maintaining reasonable performance in streaming scenarios. The method also allows specifying a language, sending all frames to the designated language expert group to avoid incorrect routing during streaming. Although we did not observe significant performance improvements, this might be more noticeable on out-of-domain test sets. Fig.2 shows the visualization of the routing results of the language router in $S11$. We found that the language router has high routing accuracy, especially for monolingual utterances. It also tends to dispatch the first few frames of English to the Zh-Group, which may correspond to frames with no speech. For CS utterances, a high routing accuracy was also observed.
Fig. 2: Visualisation of SLR’s frame-level routing results, where red represents ‘Zh’ and green represents ‘En’.
4 Conclusions
This paper presents an efficient and flexible MoE structure. Through the proposed language-group-based routing and dynamic top-$k$ mechanism, the model achieves SOTA performance, significantly improving over baselines with comparable computational costs. The study also investigates the impact of expert capacity and different top-$k$ strategies on model performance, demonstrating that language groups and the proposed dynamic top-$k$ strategy are beneficial for flexibly expanding expert capacity. In the future, we will explore the potential of the proposed model across more languages.